ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 55, 57 Koschorke, Ulrich, Universiät GH Siegen, Germany μ-invariants of link maps in arbitrary manifolds We generalize Milnor’s μ-invariants of classical links to higher dimensional link maps into arbitrary manifolds. Our approach involves the homotopy theory of configuration spaces and of wedges of spheres. In several important cases our invariants are seen to classify link maps fully up to link homotopy. Also a detailed comparison is made between μ-invariants of link maps into S^{q} × ℝ^{m−q}, 1 ≤ q < m, and into ℝ^{m}. Reference: U. Koschorke, A generalization of Milnor’s μ-invariants to higher-dimensional link maps, Topology, Vol. 36, No. 2 (1997), 301–324.